# A virtual Clebsch-Gordan formula for $\operatorname{SL}(2,\mathbb{Z})$

### v5

@@ 4. Generalities on duals and left and right comodules on $\operatorname{SL}(2,\mathbb{Z})$
You may check explicitly that the weight space decomposition of $\operatorname{S}^{2}(V)$ and $*\operatorname{S}^{2}(V)^{*}$ agree, hence we cannot distinguish these two representations using weights. You may also check explicitly there is no isomorphism of comodules $*\phi:\operatorname{S}^{2}(V)\cong\operatorname{S}^{2}(V)^{*}$. Hence the two comodules $\operatorname{S}^{2}(V)$ and $*\operatorname{S}^{2}(V)^{*}$ are not isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z})^{op}$. Similar results hold for higher symmetric powers: The comodules $\operatorname{S}^{d}(V)$ and $*\operatorname{S}^{d}(V)^{*}$ have the same weights and ranks of weight spaces, but they are not isomorphic as comodules in general.
Let $B:=A[x_{ij}](D-1)$ with $\operatorname{Spec}(B):=\operatorname{SL}(2,A)$ and let $V:=A\{e_{1},e_{2}\}$ be the following comodule on $\operatorname{SL}(2,A)$: Let $\Delta:V\rightarrow R\otimes_{A}V$ be defined by
@@
Let $\operatorname{S}^{2}(V)$ be the symmetric square of $V$ and let $*\operatorname{S}^{2}(V)^{*}$ be its dual. There is an isomorphism of comodules $*\operatorname{S}^{2}(V)\cong\operatorname{S}^{2}(V)^{*}$ iff $2$ is a unit in $B$. Hence the comodules are isomorphic when $A\cong\mathbb{Z}[\frac{1}{2}]$.
@@ 5. The second symmetric tensor of the standard comodule
In this section we prove that the second symmetric tensor $\operatorname{S}_{2}(V)$ and its dual $*\operatorname{S}_{2}(V)^{*}$ have the same weights and weight spaces but are non-isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z})$. We also prove they are isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z}[\frac{1}{2}])$. Hence a torsion free finite rank comodule on $\operatorname{SL}(2,\mathbb{Z})$ is not uniquely determined by its weights and weight space decomposition.
@@
Note: When studying group schemes and representations of group schemes one prefer to study the symmetric tensors $\operatorname{S}_{d}(V)$ over the symmetric powers $\operatorname{S}^{d}(V)$, since the symmetric tensors have better functorial properties. There is over a field $k$ of characteristic zero an isomorphism of representations $\operatorname{S}_{d}(V\otimes k)\cong\operatorname{S}^{d}(V\otimes k)$. There is an equality of classes $[\operatorname{S}_{d}(V)]=[\operatorname{S}^{d}(V)]$ in $\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))$, hence the grothendieck group $\operatorname{K}(\operatorname{SL}(2,\mathbb{Z}))$ is generated by the classes $[\operatorname{S}_{d}(V)]$ of the symmetric tensors for $d\geq 0$ an integer.
In this section we will study the symmetric tensor $\operatorname{S}_{2}(V)$ and its dual $*\operatorname{S}_{2}(V)^{*}$.
@@
$*\Delta_{2}^{*}:\operatorname{S}_{2}(V)^{*}\rightarrow R^{op}\otimes\operatorname{S}_{2}(V)^{*}$
@@
$*\tau_{2}:\operatorname{S}_{2}(V)^{*}\otimes R\rightarrow R\otimes\operatorname{S}_{2}(V)^{*}$
is the map defined by $\tau_{2}(v\otimes z):=z\otimes v$ from Lemma 4.6. Define the map $*S\otimes 1:R\otimes\operatorname{S}_{2}(V)^{*}\rightarrow R\otimes\operatorname{S}_{2}(V)^{*}$ using the involution $S$ of $R$. From Lemma 4.6 and Proposition 4.9 it follows $(\operatorname{S}_{2}(V)^{*},\Delta_{2}^{*})$ is a comodule on $R^{op}$. Let $M^{*}$ be the following matrix:
@@
One checks there is an isomorphism of comodules $*\phi:\operatorname{S}_{2}(V)\cong\operatorname{S}_{2}(V)^{*}$ iff there is a matrix $A\in\operatorname{GL}_{\mathbb{Z}}(\mathbb{Z}^{3})$ with $AM^{tr}=M^{*}A$. This gives a system of equations in the ”variables” $a_{ij}$ and one finds $\phi$ exists iff the matrix $A$ has the following form:
@@
The comodules $\operatorname{S}_{2}(V)$ and $*\operatorname{S}_{2}(V)^{*}$ have the same weights and ranks of weight spaces but are not isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z})$. The pull backs $\operatorname{S}_{2}(V)\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2]}]$ and $\operatorname{S}_{2}(V)^{*}\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]$ are isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z}[\frac{1}{2}])$.
@@
One checks the comodules have the same weights and ranks of weight spaces and an explicit calcualtion proves they are not isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z})$ using the above constructions. One also proves they are isomorphic as comodules on $\operatorname{SL}(2,\mathbb{Z}[\frac{1}{2}])$ using an explicit calculation with the matrices $M^{*}$ and $M^{tr}$. The Theorem follows. ∎
@@
Comodules on $\operatorname{SL}(2,\mathbb{Z})$ and $\operatorname{SL}(2,\mathbb{Q})$.
The comodules $*\operatorname{S}^{2}(V)^{*}$ and $*\operatorname{S}_{2}(V)^{*}$ do not have a good filtration. Hence it is easy to give explicit examples of comodules on $\operatorname{SL}(2,\mathbb{Z})$ not having a good filtration.
Let $S:=\operatorname{Spec}(\mathbb{Z})$ and $T:=\operatorname{Spec}(k)$ with $k$ the field of rational numbers. We may view the group scheme $G_{k}:=\operatorname{SL}(2,k)$ as the special fiber of the canonical map $\pi:\operatorname{SL}(2,\mathbb{Z})\rightarrow S$ and we may pull back any comodule on $\operatorname{SL}(2,\mathbb{Z})$ to $G_{k}$. Let $\tau:G_{k}\rightarrow\operatorname{SL}(2,\mathbb{Z})$ be the canonical map. Given any torsion free finite rank comodule $W$ on $\operatorname{SL}(2,\mathbb{Z})$ we get a canonical decomposition
@@
where $k^{2}$ is the standard comodule on $\operatorname{SL}(2,k)$. We would like for each integer $d\geq 0$ a classification of the comodules $W_{d}$ on $\operatorname{SL}(2,\mathbb{Z})$ with $\tau^{*}(W_{d})\cong\operatorname{S}^{d}(k^{2})$. The methods introduced above gives a method to classify all such comodules. As an example:
@@
The set of finite rank torsion free comodules on $\operatorname{SL}(2,\mathbb{Z})$ with $\tau^{*}(W)\cong\operatorname{S}^{2}(k^{2})$ are the following modules:
$*.\operatorname{S}^{2}(V),\operatorname{S}^{2}(V)^{*},\operatorname{S}_{2}(V),\operatorname{S}_{2}(V)^{*}.$

### v6

@@ 4. Generalities on duals and left and right comodules on $\operatorname{SL}_{2,\mathbb{Z}}$
You may check explicitly that the weight space decomposition of $\operatorname{S}^{2}(V)$ and $\operatorname{S}^{2}(V)^{*}$ agree, hence we cannot distinguish these two representations using weights. You may also check explicitly there is no isomorphism of comodules $\phi:\operatorname{S}^{2}(V)\cong\operatorname{S}^{2}(V)^{*}$. Hence the two comodules $\operatorname{S}^{2}(V)$ and $\operatorname{S}^{2}(V)^{*}$ are not isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}^{op}$. Similar results hold for higher symmetric powers: The comodules $\operatorname{S}^{d}(V)$ and $\operatorname{S}^{d}(V)^{*}$ have the same weights and ranks of weight spaces, but they are not isomorphic as comodules in general.
Let $B:=A[x_{ij}](D-1)$ with $\operatorname{Spec}(B):=\operatorname{SL}_{2,A}$ and let $V:=A\{e_{1},e_{2}\}$ be the following comodule on $\operatorname{SL}_{2,A}$: Let $\Delta:V\rightarrow R\otimes_{A}V$ be defined by
@@
Let $\operatorname{S}^{2}(V)$ be the symmetric square of $V$ and let $\operatorname{S}^{2}(V)^{*}$ be its dual. There is an isomorphism of comodules $\operatorname{S}^{2}(V)\cong\operatorname{S}^{2}(V)^{*}$ iff $2$ is a unit in $B$. Hence the comodules are isomorphic when $A\cong\mathbb{Z}[\frac{1}{2}]$.
@@ 5. The second symmetric tensor of the standard comodule
In this section we prove that the second symmetric tensor $\operatorname{S}_{2}(V)$ and its dual $\operatorname{S}_{2}(V)^{*}$ have the same weights and weight spaces but are non-isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. We also prove they are isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$. Hence a torsion free finite rank comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ is not uniquely determined by its weights and weight space decomposition.
@@
Note: When studying group schemes and representations of group schemes one prefer to study the symmetric tensors $\operatorname{S}_{d}(V)$ over the symmetric powers $\operatorname{S}^{d}(V)$, since the symmetric tensors have better functorial properties. There is over a field $k$ of characteristic zero an isomorphism of representations $\operatorname{S}_{d}(V\otimes k)\cong\operatorname{S}^{d}(V\otimes k)$. There is an equality of classes $[\operatorname{S}_{d}(V)]=[\operatorname{S}^{d}(V)]$ in $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$, hence the grothendieck group $\operatorname{K}(\operatorname{SL}_{2,\mathbb{Z}})$ is generated by the classes $[\operatorname{S}_{d}(V)]$ of the symmetric tensors for $d\geq 0$ an integer.
In this section we will study the symmetric tensor $\operatorname{S}_{2}(V)$ and its dual $\operatorname{S}_{2}(V)^{*}$.
@@
$\Delta_{2}^{*}:\operatorname{S}_{2}(V)^{*}\rightarrow R^{op}\otimes\operatorname{S}_{2}(V)^{*}$
@@
$\tau_{2}:\operatorname{S}_{2}(V)^{*}\otimes R\rightarrow R\otimes\operatorname{S}_{2}(V)^{*}$
is the map defined by $\tau_{2}(v\otimes z):=z\otimes v$ from Lemma 4.6. Define the map $S\otimes 1:R\otimes\operatorname{S}_{2}(V)^{*}\rightarrow R\otimes\operatorname{S}_{2}(V)^{*}$ using the involution $S$ of $R$. From Lemma 4.6 and Proposition 4.9 it follows $(\operatorname{S}_{2}(V)^{*},\Delta_{2}^{*})$ is a comodule on $R^{op}$. Let $M^{*}$ be the following matrix:
@@
One checks there is an isomorphism of comodules $\phi:\operatorname{S}_{2}(V)\cong\operatorname{S}_{2}(V)^{*}$ iff there is a matrix $A\in\operatorname{GL}_{\mathbb{Z}}(\mathbb{Z}^{3})$ with $AM^{tr}=M^{*}A$. This gives a system of equations in the ”variables” $a_{ij}$ and one finds $\phi$ exists iff the matrix $A$ has the following form:
@@
The comodules $\operatorname{S}_{2}(V)$ and $\operatorname{S}_{2}(V)^{*}$ have the same weights and ranks of weight spaces but are not isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}$. The pull backs $\operatorname{S}_{2}(V)\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2]}]$ and $\operatorname{S}_{2}(V)^{*}\otimes_{\mathbb{Z}}\mathbb{Z}[\frac{1}{2}]$ are isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$.
@@
One checks the comodules have the same weights and ranks of weight spaces and an explicit calcualtion proves they are not isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ using the above constructions. One also proves they are isomorphic as comodules on $\operatorname{SL}_{2,\mathbb{Z}[\frac{1}{2}]}$ using an explicit calculation with the matrices $M^{*}$ and $M^{tr}$. The Theorem follows. ∎
@@
Comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ and $\operatorname{SL}_{2,\mathbb{Q}}$.
The comodules $\operatorname{S}^{2}(V)^{*}$ and $\operatorname{S}_{2}(V)^{*}$ do not have a good filtration. Hence it is easy to give explicit examples of comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ not having a good filtration.
Let $S:=\operatorname{Spec}(\mathbb{Z})$ and $T:=\operatorname{Spec}(k)$ with $k$ the field of rational numbers. We may view the group scheme $G_{k}:=\operatorname{SL}(2,k)$ as the special fiber of the canonical map $\pi:\operatorname{SL}_{2,\mathbb{Z}}\rightarrow S$ and we may pull back any comodule on $\operatorname{SL}_{2,\mathbb{Z}}$ to $G_{k}$. Let $\tau:G_{k}\rightarrow\operatorname{SL}_{2,\mathbb{Z}}$ be the canonical map. Given any torsion free finite rank comodule $W$ on $\operatorname{SL}_{2,\mathbb{Z}}$ we get a canonical decomposition
@@
where $k^{2}$ is the standard comodule on $\operatorname{SL}(2,k)$. We would like for each integer $d\geq 0$ a classification of the comodules $W_{d}$ on $\operatorname{SL}_{2,\mathbb{Z}}$ with $\tau^{*}(W_{d})\cong\operatorname{S}^{d}(k^{2})$. The methods introduced above gives a method to classify all such comodules. As an example:
@@
The set of finite rank torsion free comodules on $\operatorname{SL}_{2,\mathbb{Z}}$ with $\tau^{*}(W)\cong\operatorname{S}^{2}(k^{2})$ are the following modules:
$\operatorname{S}^{2}(V),\operatorname{S}^{2}(V)^{*},\operatorname{S}_{2}(V),\operatorname{S}_{2}(V)^{*}.$